I am a mom trying to remember the steps to this problem to help my son.
4 x (9+3)= (4 x n) + (4 x 3)

Answer by nyc_function(2741)

Do not use the letter x to represent multiplication in algebra because x is

used a lot as a variable and will produce confusion. In place of x, use * or

parentheses, which also represent multiplication.

We are looking for n.

This equation is called a linear equation.

4 x (9+3)= (4 x n) + (4 x 3)

4 * 12 = 4*n + 12

48 = 4n + 12

48 - 12 = 4n

36 = 4n

To find the value of n simply divide 36 by 4 and we are done.

36/4 = n

9 = n

The answer is 9.

=================
Here is the proof:

In the original equation given, let n = 9. If we get the same answer on both

sides of the equation, then we know for sure that n = 9 in this sample.

Let n = 9

4 x (9+3)= (4 x n) + (4 x 3)

4 x (9+3)= (4 x 9) + (4 x 3)

4 * 12 = 36 + 12

48 = 48....It checks!!

We now know for sure that n = 9.

By former substitute teacher.